OpenStudy (anonymous):

Solve each equation 2x-y=5 3x+y=5

OpenStudy (anonymous):

If you use the elimination method in this, you can get x alone. add the two equations together. 2x-y+3x+y=5+5 simplify 5x=10 Divide by 5 to get x alone, then plug x back into one of the equations to find y... When you're done post what you get and I'll check if you're right.

OpenStudy (anonymous):

Solve by graphing

OpenStudy (anonymous):

Then this is easy to do yourself... I don't have a graph to show you. Just make one. Do one graph at a time. I suggest inputting 0,2,4,6 for x.